OpenStudy (anonymous):

For each linear function of phi on V, find u such that phi(v)= for every v is an element of V

OpenStudy (anonymous):

1) \[\phi: R^3\rightarrow R \] defined by \[\phi (x,y,z)= x+2y-3z\] 2) \[\phi: C^3\rightarrow C \] defined by \[\phi (x,y,z)= ix+(2+3i)y+(1-2i)z\]

OpenStudy (anonymous):

Isn't it clear ? U in first case is U=(1,2,-3) In second \[(i, (2+3i), (1-2i))\]

OpenStudy (anonymous):

I was right. Thanks!